I have a 3000 gallon glycol heating system. I need to achieve a 50% mixture of glycol and water. Right now my 3000 gallons is around a 30% mixture, I want to buy 100% glycol and mix it to bring the 3000 gallons back up to a 50% mixture. Im trying to find out how many gallons i need to do this. I know it will raise the 3000 gallons as well which means I wont need the entire 3000 of the 30%. I couldnt begin to try and wrap my head around this. Anything you could do to help would be greatly appreciated.
Answer by Alan3354(69443)

I have a 3000 gallon glycol heating system. I need to achieve a 50% mixture of glycol and water. Right now my 3000 gallons is around a 30% mixture, I want to buy 100% glycol and mix it to bring the 3000 gallons back up to a 50% mixture. Im trying to find out how many gallons i need to do this. I know it will raise the 3000 gallons as well which means I wont need the entire 3000 of the 30%. I couldnt begin to try and wrap my head around this. Anything you could do to help would be greatly appreciated.
======================
Drain x gallons of the 30% to make space for the 100% to be added.
----
(3000 - x)*30 + x*100 = 3000*50
x = 6000/7 gallons to drain
x =~ 857 gallons to drain & replace.